ADNAN MUHAMMED

Grade 12,

A uniform solid sphere of relative density 5 is released in water filled in a long vertical tube. Its terminal velocity acheived is V. If another uniform solid sphere of same material but double the radius is released in the same water then the terminal velocity achieved will be.

A uniform solid sphere of relative density 5 is released in water filled in a long vertical tube. Its terminal
velocity acheived is V. If another uniform solid sphere of same material but double the radius is released
in the same water then the terminal velocity achieved will be.

Grade:12th Pass

1 Answers

vikas askiitian expert
13 years ago

force (F) = 6piRkV                 (V is terminal velocity)

Fi/Ff = (Ri/Rf).(Vi/Vf)   ..........1              (i , f denotes initial & final)

 

F = vg(p - pw)             (v is volume)

F1/F2 = v1/v2                        (v = 4/3 piR3 )

F1/F2 = (Ri/Rf)2       ...............2

equating 1 & 2

(Ri/Rf)2 = (Vi/Vf)

final radius is double so

 Vi/Vf = 1/4

 Vf = 4V

approve if u like my ans